# Towards AI-Assisted Synthesis of Verified Dafny Methods

### v1

@@ 3.6. Temperature Tuning
The “temperature” is an important parameter of all LLMs that controls the randomness of the decoding process (OpenAI, 2023a; Rohan Anil et al., 2023). In doing experiments with LLMs, it is important to identify the best temperature for each experiment. Therefore, before running our experiments in the entire MBPP-san-DFY dataset, we first performed a temperature tuning experiment. We chose a random sample of 50 problems from our test dataset MBPP-san-DFY of 178 problems employing the sampling procedures with 90% confidence interval and 10% margin of error. Next, for each type of prompt, we executed these 50 problems in four different temperatures $T\in\{0.0,0.25,0.5,0.75\}$ in both GPT-4 and PaLM-2 . We evaluated the synthesized Dafny methods with our evaluation metrics verify@k where $k\in\{1,3,5\}$. As shown in Figure 5 , we observed that in Contextless and Signature prompt, GPT-4 shows best performance in T=0.75. However, in Dynamic Few-Shot prompt, GPT-4 is most effective in T=0.5. PaLM-2 demonstrates better performance with T=0.0 in Contextless prompt. And in Signature and Dynamic Few-Shot Prompts PaLM-2 works good in T=0.5. Thus, we used those values for GPT-4 and PaLM-2 , and their respective prompts.
Figure 5. Temperature tuning for GPT-4 and PaLM-2 on a sample of 50 problems in four different temperatures $T\in\{0.0,0.25,0.5,0.75\}$ for three different prompts. The values in vertical axis represents the verified program synthesis rate in verify@k (k=5).
@@ 3.7. Experimental Setup
We developed scripts to execute different prompts on different temperatures through Open-AI’s API for GPT-4 and Google’s API for PaLM-2. For verifying the code, we employed the latest Dafny 4.0.0 version  [^5] https://dafny.org/blog/2023/03/03/dafny-4-released/. In March 3, 2023, a new major version, Dafny 4.0.0, has been released that broke backwards compatibility. Both GPT-4 and PaLM-2 may not know anything about this latest version. In our prompts and few-shot examples we did not specify which version of Dafny should be used. We verified our MBPP-DFY-50 dataset in Dafny 4.0.0. For executing the experiments and verifying synthesized Dafny methods, we used a MacBook Pro (13-inch, 2020), having 2 GHz Quad-Core Intel® Core™ i5 processor, 16GB RAM, 512 GB SSD and running macOS Ventura 13.4.
@@ 4. Results and Analysis
To evaluate LLM synthesis of Dafny methods, we queried all 178 problems in three prompts with their tuned temperatures (described in Section 3.6). However, we noticed that LLMs cannot generate any Danfy method snippets if they misunderstood the problem or had insufficient knowledge to solve those problems. In such cases, LLMs returned a textual explanation why it cannot generate any Dafny method snippets: we filtered out those responses. We then analyzed the verification results quantitatively and qualitatively.
@@ 4.1.1. Verification:
Table 1 summarizes the overall verification results for three prompts queried with their tuned temperatures. For example, in Contextless prompting at k=5, GPT-4  synthesized 104 verified methods out of 178 problems, achieving a verification success rate 58.42%. We notice that the verification success rate improves with the increasing number of attempts. For instance, in Contextless for GPT-4 in k=1, the verification success rate is 32.58% whereas at k=3 and k=5 it jumps to 53.37% and 58.42% respectively. PaLM-2 was unable to generate a single verified method for Contextless prompting even with several attempts. In Signature prompting, the verification success rate also improves with the number of attempts. Here, GPT-4  generated 95 Dafny verified methods (53.37%), which is slightly lower than Contextless prompting. PaLM-2   gets a boost at k=5, synthesizing 12 (6.74%) verified methods. Both GPT-4 and PaLM-2 attain the best performance in Dynamic Few-shot prompting. GPT-4 got 114 verified methods (64.04%) in k=5 that surpass Contextless prompting’s performance. PaLM-2 achieved significant improvement with a 29.21% verification success which is around 4X performance boost compared to Signature prompting.
Table 1. Summary of verified Dafny method synthesis on 178 problems. For each prompt, LLMs responses were generated with their tuned temperature (T) in verify@k (where $k\in\{1,3,5\}$)
@@ 4.1.2. Postconditions:
In a naive interpretation, these results could indicate that the complicated third prompt is not worth it, at least for GPT-4 . However, that is far from being the case. Verification success does not mean that the solution was properly formulated. Postconditions, specifically, are essential in formally specifying what methods are supposed to do. The absence of postconditions is a strong sign that the solution is not usefully verifiable: methods with no specifications trivially avoid being wrong.
Table 2. Summary of specifications synthesis result, in both verified and unverified methods. For each prompt, LLMs responses were generated with their tuned temperatures at verify@k (k=5).
@@
For both the verified and unverified methods, we looked into the formal specification synthesis results. In Contextless prompting as shown in Table 2, we noticed that, although, GPT-4  generated 104 verified methods, only 56 of them contain postconditions. Since the remaining 48 verified methods have no postconditions, indeed, they cannot be considered formally verified. Regarding unverified methods, GPT-4 generates postconditions in 65.75% of them. Similar to GPT-4 , PaLM-2 also shows better effectiveness (59.09%) in synthesizing postconditions in unverified methods. However, since these methods are unverified, it may happen that either these postconditions are wrong or the syntactic or semantic implementation of these methods is incorrect to prove the correct postconditions. In our qualitative analyses we manually examine these methods and their specifications to identify the actual reasons (see Section 4.2).
In Signature prompting, the presence of postconditions gets worse. Only 30 out of 95 synthesis methods from GPT-4 have postconditions whereas for PaLM-2 only 1 method out of 12 contains postconditions. For unverified methods the presence of postconditions drastically drops for both GPT-4 (31.57%) and PaLM-2 (38.27%). It indicates that including additional contexts in the prompt did not bring any substantial benefit.
However, in Dynamic Few-Shot prompting, the specification synthesis results outstandingly improves for both GPT-4 and PaLM-2 . According to Table 2, 100% of both verified and unverified methods have postconditions. The primary reason for having postconditions in all synthesized methods is the few-shot examples with the CoT process. We designed this prompt to dynamically select five few-shot example methods based on their postconditions’ semantic similarity, and each of these examples includes a CoT process describing how to determine the post-conditions in natural instructions and then in Dafny method in the implementation. The design of this prompt made the LLMs follow the manner by which the examples were implemented.
4.1.3. Preconditions:
In a Dafny method, preconditions are only required when certain properties or assumptions must be hold before invoking the method. Having preconditions is not a mandatory formal specification that needs to be proved by the methods’ implementation.
In our experiment with Contextless prompting, we found that 61.53% of GPT-4 generated verified methods contain preconditions and 39.42% contain both pre-and postconditions (Table 2). Like postconditinos, the presence of preconditions drops in Signature prompting where only 48.42% of verified methods contain preconditions and 22.10% have both pre-and postconditions. Unfortunately, in Contextless and Signature prompting, PaLM-2 specification synthesis performance is trivial. However, for both GPT-4 and PaLM-2 , Dynamic Few-Shot prompting enhances the specification synthesis performance. As demonstrated in Table 2, on average 64% of verified and unverified programs for both GPT-4 and PaLM-2 contain pre-and postconditions. It substantiate that the design of Dynamic Few-Shot prompt considerably strengthens LLMs specification synthesis performance.
@@ 4.1.4. Invariants:
When a Dafny method has a loop, it should contain invariants that need to be maintained in each iteration of the loop. As shown in Table 2, in Contextless prompt, on an average 38% GPT-4 synthesis methods contain invariants. In Signature prompting, the number increases slightly for verified methods, however, the presence of invariants exceptionally enhances for Dynamic Few-Shot prompting in GPT-4 synthesis methods. According to Table 2, 50% of verified and 93.44% of unverified methods contain at least one loop invariant. On the contrary, PaLM-2 struggles to synthesize loop invariants for both verified and unverified methods in Contextless and Signature prompting. Only the Dynamic Few-Shot prompting outstandingly improves its performance to have loop invariants in 25% of verified and 73.17% for unverified methods respectively.
@@ 4.2. Qualitative Analysis
This section describes our manual investigation. The first two authors manually examined all prompt responses, assessing the quality of the formal specifications, specifically the postconditions of the verified Dafny methods – again, postconditions are the key element for a correct formal specification. During this manual inspection, when we have any doubt about the implementation or specifications, we looked into our ground-truth dataset, MBPP-san, and discussed with each other. The two authors independently performed this manual tagging and reached a high Inter-Rater Reliability (IRR) (Hallgren, 2012)(93.91% Cohen Kappa (Viera et al., 2005)), which ended up with consensus at the end. We summarize our qualitative analysis findings in Table 3. The complete analysis results are available in our companion repository 8, including all prompts, LLMs’ responses, verification results, Dafny method snippets with our manual tagging.
Table 3. Summary of qualitative analysis of specifications synthesis in verified methods. For each prompt, LLMs responses were generated with their tuned temperatures at verify@k (k=5).
@@
4.2.1. Postconditions:
The goal of a postcondition is to make sure that the implementation is correct based on the proofs. Here the term ”correct” implies that the postcondition is strong enough to precisely capture all possible output values. We tagged the quality of postconditions into three groups, Strong, Weak and Wrong. In Signature prompting, PaLM-2 got only one verified method and that has Strong postconditions. With regard to, Dynamic Few-Shot prompt, we found that in 14 cases, PaLM-2 misunderstood problems’ description and generated wrong method implementation for those problems. However, it got Strong postconditions in 35 methods and Weak postconditions in 3 methods only. Although GPT-4 got 56 verified programs in Contextless prompting, we encountered 15 of them have Weak postconditions. Table 4.2.1 exhibits such an example of Weak postconditions. On the other hand in Signature prompting 9 of verified methods contain Weak postconditions which are very trivial specifications to verify correct implementation.
In Dynamic Few-Shot, GPT-4 ’s specification synthesis result is impressive. Only 3 of verified methods got Weak postconditions and in 8 cases it misunderstood the problems’ description generated wrong methods. One interesting example of such cases where GPT-4 misunderstood the problem description and generated Wrong method is illustrated in Table 4.2.1. Specifically in Dynamic Few-Shot, GPT-4 got 90.35%(103/114) of verified methods containing Strong postconditions.
Table 4. Examples of LLMs synthesis Strong, Weak and Wrong postconditions; Required and Optional preconditions; invariants subsuming postconditions, and modularization of proofs.
@@ 4.2.1. Postconditions:
/*Write a method in Dafny to interleave 3 sequence of the same
length into a single flat sequence.*/
method SplitAndAppend(l: seq<int>, n: int) returns (r: seq<int>)
    requires n >= 0 && n < |l|
    ensures |r| == |l|
    ensures forall i :: 0 <= i < |l| ==> r[i] == l[(i + n) % |l|]
@@
function Power(base: int, exponent: int): int //.. PROOF MODULARIZATION

### v2

@@ 3.6. Temperature Tuning
The “temperature” is an important parameter of all LLMs that controls the randomness of the decoding process (OpenAI, 2023a; Rohan Anil et al., 2023). In doing experiments with LLMs, it is important to identify the best temperature for each experiment. Therefore, before running our experiments on the entire MBPP-san-DFY dataset, we first performed a temperature tuning experiment. We chose a random sample of 50 problems from our test dataset MBPP-san-DFY of 178 problems employing the sampling procedures with 90% confidence interval and 10% margin of error. Next, for each type of prompt, we executed these 50 problems at four different temperatures $T\in\{0.0,0.25,0.5,0.75\}$ in both GPT-4 and PaLM-2 . We evaluated the synthesized Dafny methods with our evaluation metrics verify@k where $k\in\{1,3,5\}$. Figure 5 shows that in Contextless and Signature prompts, GPT-4 performs best at T=0.75, although with Dynamic Few-Shot prompt, GPT-4 is most effective at T=0.5. PaLM-2 demonstrates better performance with T=0.0 for Contextless prompt, but in Signature and Dynamic Few-Shot prompts PaLM-2 works better at T=0.5.
Figure 5. Temperature tuning for GPT-4 and PaLM-2 on a sample of 50 problems at four different temperatures $T\in\{0.0,0.25,0.5,0.75\}$ for three different prompts. The vertical axis shows the verified Dafny method synthesis rate in verify@k (k=5).
@@ 3.7. Experimental Setup
We developed scripts to execute prompts at different temperatures through OpenAI’s API for GPT-4 and Google’s API for PaLM-2. For verifying the code, we employed the latest Dafny 4.0.0 version  [^5] https://dafny.org/blog/2023/03/03/dafny-4-released/. Dafny 4.0.0, released in March 2023, broke backwards compatibility, so it is likely neither GPT-4 nor PaLM-2 would have this latest version in their training data. In our prompts and few-shot examples we did not specify which version of Dafny should be used. For executing the experiments and verifying synthesized Dafny methods, we used a MacBook Pro (13-inch, 2020), having 2 GHz Quad-Core Intel® Core™ i5 processor, 16GB RAM, 512 GB SSD and running macOS Ventura 13.4.
@@ 4. Results and Analysis
To evaluate LLM synthesis of Dafny methods, we queried all 178 problems in three prompts with their tuned temperatures (as described in Section 3.6). We noticed that LLMs could not generate any Dafny methods if they were unable to generate the correct postconditions or if they had insufficient context data to solve those problems. In such cases the LLMs returned a textual explanation why they could not generate any Dafny code; we filtered out those responses. We then analyzed the verification results quantitatively and qualitatively.
@@ 4.1.1. Verification:
Table 1 summarizes the overall verification results for three prompts queried with their tuned temperatures. For example, in Contextless prompting at k=5, GPT-4  synthesized 104 verified methods out of 178 problems, achieving a verification success rate 58.42%. We notice that the verification success rate improves with the increasing number of attempts. For instance, in Contextless prompt at k=1, GPT-4 ’s verification success rate is 32.58% whereas at k=3 and k=5 it jumps to 53.37% and 58.42% respectively. PaLM-2 was unable to generate a single verified method for Contextless prompting even with several attempts. In Signature prompting, the verification success rate also improves with the number of attempts. Here, GPT-4  generated 95 verified Dafny methods (53.37%), which is slightly lower than Contextless prompting. PaLM-2   gets a boost at k=5, synthesizing 12 (6.74%) verified methods. Both GPT-4 and PaLM-2 attain the best performance in Dynamic Few-shot prompting. GPT-4 verified 114 methods (64.04%) at k=5 that surpasses Contextless prompting’s performance. PaLM-2 achieved significant improvement with a 29.21% verification success which is around four times of its performance with Signature prompting.
Table 1. Summary of verified Dafny method synthesis on 178 problems. For each prompt, LLMs responses were generated with their tuned temperature (T) at verify@k (where $k\in\{1,3,5\}$)
@@ 4.1.2. Postconditions:
In a naive interpretation, these results could indicate that the complicated RAG prompt is not worth it, at least for GPT-4 ; however, that is not the case. Verification success does not mean that the solution was properly formulated. Postconditions are essential in formally specifying what methods are supposed to do. The absence of postconditions is a strong sign that the solution is not usefully verified: methods with no specifications trivially avoid being wrong.
For both the verified and unverified methods, we looked into the formal specification synthesis results. In Contextless prompting as shown in Table 2, we noticed that, although GPT-4  generated 104 verified methods, only 56 of them contain postconditions. Since the other 48 verified methods have no postconditions, they cannot be considered formally verified. Regarding unverified methods, GPT-4 generates postconditions for 65.75% of them. Similar to GPT-4 , PaLM-2 also shows better effectiveness (59.09%) in synthesizing postconditions in unverified methods. Since these methods are unverified, however, it may happen that either the postconditions are wrong, that the syntactic or semantic implementation of these methods are wrong, or that both specifications and method implementations are correct, but that the Dafny verifier was unable to prove the method implementation in fact maintained the postconditions. In our qualitative analyses we manually examine these methods and their specifications to identify the actual reasons (see Section 4.2).
In Signature prompting, the presence of postconditions gets worse. Only 30 out of 95 synthesis methods from GPT-4 have postconditions whereas for PaLM-2 only 1 method out of 12 contains postconditions. For unverified methods the presence of postconditions drastically drops for both GPT-4 (31.57%) and PaLM-2 (38.27%), indicating that any additional contexts in the prompt did not bring any substantial benefit.
With Dynamic Few-Shot RAG prompting, however, the specification synthesis results improve significantly for both GPT-4 and PaLM-2 . As shown in Table 2, 100% of both verified and unverified methods have postconditions. The primary reason for having postconditions in all synthesized methods is the few-shot examples with the CoT process. We designed this prompt to dynamically select five methods based on their postconditions’ semantic similarity, and each of these examples includes a CoT process describing how to determine the postconditions in natural language, and then in Dafny in the implementation. The design of this prompt makes the LLMs follow the manner by which the examples are presented.
4.1.3. Preconditions:
In a Dafny method, preconditions are only required when certain properties or assumptions must hold before invoking the method. For example, if a method reads an array element, preconditions can be used to help ensure that the method does not attempt to access array elements that are out-of-bounds.
In our experiment with Contextless prompting, we found that 61.53% of GPT-4 generated verified methods contain preconditions and 39.42% contain both pre-and postconditions (Table 2). Like postconditions, the presence of preconditions drops in Signature prompting where only 48.42% of verified methods contain preconditions and 22.10% have both pre-and postconditions. Unfortunately, in Contextless and Signature prompting, PaLM-2 specification synthesis performance is trivial. For both GPT-4 and PaLM-2 , however, Dynamic Few-Shot RAG prompting enhances the specification synthesis performance. As shown in Table 2, on average 64% of verified and unverified programs for both GPT-4 and PaLM-2 contain pre-and postconditions. Once again, Dynamic Few-Shot prompts perform significantly better than Contextless and Signature prompts.
Table 2. Summary of specifications synthesis result, in both verified and unverified methods. The number in the denominators represents the total number of verified and unverified methods respectively. For each prompt, LLMs responses were generated with their tuned temperatures at verify@k (k=5).
@@ 4.1.4. Invariants:
When a Dafny method has a loop, it should contain invariants that need to be maintained in each iteration of the loop. As shown in Table 2, in Contextless prompting, an average 38% GPT-4 synthesized methods contain invariants. In Signature prompting, the number increases slightly for verified methods, however, in Dynamic Few-Shot prompting, GPT-4 does far better at generating invariants. As shown in Table 2, 50% of verified and 93.44% of unverified methods contain at least one loop invariant. On the other hand, PaLM-2 struggles to synthesize loop invariants for both verified and unverified methods in Contextless and Signature prompting. Again, Dynamic Few-Shot prompting improves its performance to have loop invariants in 25% of verified and 73.17% for unverified methods respectively.
@@ 4.2. Qualitative Analysis
This section describes our manual investigation of LLMs’ Dafny synthesis performance. The first two authors manually examined all prompt responses, assessing the quality of the formal specifications, specifically the postconditions of the verified Dafny methods – again, postconditions are the key element for a correct formal specification. During this manual inspection, when we have any doubt about the implementation or specifications, we looked into our ground-truth dataset, MBPP-san, and discussed with each other. The two authors independently performed this manual tagging and reached a high Inter-Rater Reliability (IRR) (Hallgren, 2012)(93.91% Cohen Kappa (Viera et al., 2005)), which ended up with consensus at the end. We summarize our qualitative analysis findings in Table 3. The complete analysis results are available in our companion repository (Section 8), including all prompts, LLMs’ responses, verification results, and Dafny method snippets with our manual tagging.
4.2.1. Postconditions:
The goal of a postcondition is to make sure that the implementation is correct based on the proofs. Here the term ”correct” implies that the postcondition is strong enough to precisely capture all possible output values. A postcondition can be weak or less restrictive that allows a boarder range of possible outcomes. In some cases, a postcondition might be incorrect to accurately capture the intended outcomes and making it difficult or impossible to prove.
We tagged the quality of postconditions into three groups, Strong, Weak and Wrong. In Signature prompting, PaLM-2 had only one verified method and that has Strong postconditions. With regard to, Dynamic Few-Shot prompt, we found that in 14 cases, PaLM-2 misunderstood problems’ description and generated wrong method implementation for those problems. However, it generated Strong postconditions in 35 methods and Weak postconditions in 3 methods only. Although GPT-4 got 56 verified programs in Contextless prompting, we encountered 15 of them have Weak postconditions. Table 4.2.1 exhibits such an example of Weak postconditions. On the other hand in Signature prompting 9 of verified methods contain Weak postconditions which are very trivial specifications to verify correct implementation.
In Dynamic Few-Shot, GPT-4 ’s specification synthesis result is impressive. Only 3 of verified methods had Weak postconditions and in 8 cases it misunderstood the problems’ description, generating wrong methods. One interesting example of such cases is illustrated in Table 4.2.1. Specifically in Dynamic Few-Shot RAG prompt, 90.35% (103/114) of verified methods synthesized from GPT-4 contain Strong postconditions. Considering the overall dataset, 58% of methods (103/178) had strong (i.e. correct) specifications that successfully verified.
Table 3. Summary of qualitative analysis of specifications synthesis in verified methods. The number in the denominator represents the total number of verified methods containing specifications. For each prompt, LLM responses were generated with their tuned temperatures at verify@k (k=5).
@@ 4.2.1. Postconditions:
Table 4. Examples of LLMs synthesis Strong, Weak and Wrong postconditions; Required and Optional preconditions; invariants subsuming postconditions, and auxiliary definitions.
@@
/*Write a method in Dafny to interleave 3 sequence of the same length into
a single flat sequence.*/
method InterleaveThreeSequences(seq1: seq<int>, seq2: seq<int>, seq3: seq<int>)
returns (interleavedSeq: seq<int>)
requires |seq1| == |seq2| == |seq3|
ensures |interleavedSeq| == 3 * |seq1|
//strong postcondition will also check the relative order of the sequence values.
@@
function Power(base: int, exponent: int): int //.. Auxiliary Definitions
@@
4.2.2. Preconditions:
A precondition is an annotation that must be true before a Dafny method is invoked. When a method is called, it is caller’s responsibility to make sure that the preconditions are satisfied. The caller of the method then gets to assume that the postconditions hold after the method returns  (Dafny, 2023). To keep in mind that, in our analysis, we did not call LLMs synthesis methods. In our manual investigation, we only tagged preconditions into two groups, (i) Required, which is necessary to prove the postconditions and implementation; (ii) Optional, its presence is not required for verification. Table 3 demonstrates the summary of our preconditions analysis. Examples of LLMs synthesis Required and Optional preconditions are illustrated in Table 4.2.1.
4.2.3. Invariants:
We reviewed loop invariants in all verified methods that have postconditions. To determine loop invariants in Dafny, a common trick is to work backwards from the postconditions and identify the loop guard (Dafny, 2023). While examining responses, we recognized the same patterns followed by the LLMs. We observed that GPT-4 and PaLM-2 both determined loop invariants mostly based on the postconditions and in most of the problems loops invariants subsume some or whole part of the method’s postconditions as shown in Table 4.2.1. We found that all synthesized verified methods in three different prompts containing Strong postconditions, also have correct loop invariants. And when the postconditions were Weak or Wrong, loop invariants followed the postconditions. In many cases we also noticed that GPT-4 guaranteed loop termination with the explicit usages of decreases statement on loop guards.
4.2.4. Auxiliary Definitions:
Dafny, supports predicate and function to write modularized and readable code (Dafny, 2023). When we reviewed all verified synthesized methods, auxiliary definitions were only generated in response to Dynamic Few-Shot prompts. We found that only 4 out of 52 PaLM-2 synthesized verified methods had auxiliary definitions whereas 20 out of 103 GPT-4 verified methods have the usages of predicate or function. Table 4.2.1 shows an example of such auxiliary definitions with the usages of function.
4.3. Error Analysis
Table 5. Summary of different types of errors found in unverified methods. The number in the denominators represents the total number of unverified methods. For each prompt, LLMs responses were generated with their tuned temperatures at verify@k (k=5).
Error Types Error Name GPT-4 PaLM-2
Prompts Prompts
Contextless Signature
Dynamic
Few-Shot
Contextless Signature
Dynamic
Few-Shot
Compilation Errors parse errors 51/73 44/80 16/61 176/176 152/162 43/124
resolution/type errors 14/73 30/80 15/61 0/176 10/162 66/124
Verification Errors loop invariant violation 4/73 1/80 23/61 0/176 0/162 10/124
postcondition might not hold 3/73 1/80 13/61 0/176 0/162 7/124
bound errors 0/73 1/80 2/61 0/176 0/162 5/124
index out of range 2/73 2/80 5/61 0/176 0/162 2/124
possible division by zero 0/73 0/80 0/61 0/176 0/162 1/124
verification time out 1/73 0/80 0/61 0/176 0/162 0/124
We manually analyzed the error logs for all unverified methods and identified two major types of errors. We then examined the error types and classified them into several subgroups. Table 5 demonstrates the presence of different types of errors appeared in all unverified methods.
4.3.1. Compilation Errors:
We found compilation errors when a Dafny method was not verified because of incorrect Dafny syntax or unknown types. These errors appeared in two forms (i) parse errors, having wrong Dafny syntax and (ii) resolution/type errors, when a method implementation contains usages of any unknown types. In our manual inspection, we observed that the majority of LLMs synthesis methods were not verified because of compilation errors. As shown in Table 5, in Contextless and Signature prompting, 100% and 93.83% of PaLM-2 synthesis methods were unverified because of parse errors. In Dynamic Few-Shot RAG prompting, parse errors were minimized, nevertheless, resolution/type errors appeared in around 53.22% of cases. We inspected the reasons for resolution/type errors and found that PaLM-2 generated Dafny methods assuming the presence of basic library methods/functions without providing any implementation. Consequently, the verifier raised a resolution/type errors because of the usages of unknown method/function. Similarly, GPT-4 also struggled generating correct Dafny syntax in Contextless and Signature prompting. In Contextless, 51 out of 73 unverified methods have parse errors. It gradually overcome and improved its performance in Dynamic Few-Shot prompting where only 16 out of 61 unverified methods had parse errors. Regarding, resolution/type errors, GPT-4 generated such errors most (30/80) in Signature prompting whereas in Dynamic Few-Shot prompting, GPT-4 had only 15 of such errors out of 61 unverified methods.
4.3.2. Verification Errors:
These types of errors occurred when a Dafny method has correct syntax but the specifications or implementation were incorrect. We found six different types of verification errors, shown in Table 5. Most of the syntactically correct unverified methods failed to verify because of loop invariant violation and/or postcondition might not hold errors. In Contextless and Signature prompting, there were very few verification errors since most of the methods had compilation errors. However, in Dynamic Few-Shot prompting, GPT-4 had loop invariant violation in 23 cases and postcondition might not hold in 13 out of 61 unverified methods. We observed that these errors appeared for two reasons: (i) the specifications (postconditions or invariants) were correct but the implementation was incorrect to prove the specifications; and (ii) the implementation was correct but the specifications were incorrect.
4.4. Summary of Findings
RQ1 [Contextless Prompting]: Our experiments show that, given a simple problem description, LLMs can generate plausible Dafny methods, although the resulting methods are not necessarily syntactically correct. Specifically, GPT-4 is capable of generating Dafny methods, but often cannot generate the formal specifications. Without proper formal specifications, the internal verification annotations, when they exist, become moot.
RQ2 [Signature Prompting]: Additional context — method signature and test cases — reduced the performance of both the LLMs that we evaluated. This contrasts with LLMs undertaking more classical code synthesis tasks, where additional contextual information generally improves their performance.
RQ3 [Dynamic Few-Shot Prompting]: Our results show that Chain of Thought (CoT) prompts, separately prompting for specifications, and then chaining those generated specifications in a second prompt with retrieval augmentation generated (i.e. RAG) semantically similar few-shot examples, significantly increased performance. Engineering these prompts led both LLMs to generate explicit pre- and post-conditions for all the test problems. Specifically for GPT-4 , 100% of all generated methods contained formal specifications, and 58% of them (103/178) had strong (i.e. correct) specifications that successfully verified.
5. Related Work
5.1. Program Synthesis and Verification with Dafny
In the last 20 years, formal methods for software synthesis (Gulwani et al., 2017) and verification (Ringer et al., 2019) have moved from an esoteric research topic (Leino and Nelson, 1998) to a set of increasingly practical tools, and from doctoral study to undergraduate degrees (Jones and Misra, 2021). Tools such as Dafny, SAW, and SPIN are increasingly mature enough to support industrial application (Greengard, 2021; Wayne, 2018): indeed a main barrier to adoption remains a lack of software engineers trained in their use (Garavel et al., 2020).
Dafny is used in production to develop verified implementations encryption algorithms (Yang et al., 2023), Ethereum Virtual Machine (EVM) bytecodes (Cassez et al., 2023), and even quantum circuitry (Li et al., 2022b). Dafny is also the topic of ongoing research in verification and synthesis. Irfan et al. (2022) introduced XDsmith, a fuzzing differential testing framework that generates random Dafny programs with known verification output. XDsmith conducts differential testing on the Dafny compiler to evaluate Dafny’s soundness and precision. In the Amazon Automated Reasoning group, Chakarov et al. (2022) developed an approach to generate counterexamples in Dafny syntax when the Dafny verifier cannot establish a valid proof.
This work testifies to the ongoing relevance of Dafny-style program verification in research, and increasingly in practice. Our results indicate that LLMs are likely to be a fruitful direction for assisting with Dafny programming, in both research and practice.
5.2. LLMs for Formal Methods
Researchers have been actively working on developing automated proof synthesis and theorem provers (Sanchez-Stern et al., 2020). Several efficient proof synthesis tools have been developed based on pre-computed facts reasoning, heuristic search and identifiers, such as CoqHammer (Czajka and Kaliszyk, 2018), TacTok (First et al., 2020), Passport (Sanchez-Stern et al., 2023). Recently developed LLMs based approaches, for instance, ASTactic (Yang and Deng, 2019), Diva (First and Brun, 2022) Thor (Jiang et al., 2022a) and Baldur (First et al., 2023) outperform the previous approaches to synthesize whole or partial proofs and automatically prove theorems employing the interactive proof assistants Coq and Isabelle/HOL.
Apart from proof synthesis and theorem proving, LLMs have also been applied to translate natural language into formal specifications, proofs, and theorems (Qiao et al., 2023; Meadows and Freitas, 2022; Li et al., 2023). Wu et al. translated natural language problems into Isabelle/HOL (Wu et al., 2022), while Cunningham et al. experimented generating proofs for code verification (Cunningham et al., 2022). To enhance reasoning, Madaan et al. generated event graphs (Madaan et al., 2022) and found that LLMs trained on source code outperformed those trained only on text.
Most contemporary LLMs have limitations when it comes to algorithmic reasoning (Austin et al., 2021) although this may be mitigated by careful prompting (Zhou et al., 2022). While GPT-4 can generate plausible texts about mathematical problems, it cannot currently deliver correct mathematical proofs or solutions (Frieder et al., 2023). The Minerva model, however, is pre-trained on a mathematics corpus (First et al., 2023), and also aims to generate proofs in Isabelle/HOL.
Learning techniques are in vogue in software engineering research (Watson et al., 2022; Yang et al., 2022; Allamanis et al., 2018). LLMs have been evaluated across a range of tasks such as code generation, code completion, code summarization (Tufano et al., 2023; Nashid et al., 2023), code repair (Yasunaga and Liang, 2020; Sintaha et al., 2023; Mashhadi and Hemmati, 2021), program transformation and synthesis (Austin et al., 2021), test generation (Mastropaolo et al., 2023; Tufano et al., 2022a), bug fixing (Chakraborty and Ray, 2021; Tufano et al., 2019), code review (Tufano et al., 2022b; Li et al., 2022a) — for some tasks, performing as well as or better than the state of the art (Watson et al., 2022; Yang et al., 2022; Allamanis et al., 2018), especially when prompts are carefully constructed (Nashid et al., 2023). Recent studies suggest that applying LLMs to code-related tasks has the potential to improve developer productivity (Nashid et al., 2023; Chakraborty and Ray, 2021; Tufano et al., 2019).
All this work (both formal methods and programming tasks) focuses on careful training of LLMs in mathematics — or otherwise extending their reasoning power. Our results explore an alternative approach: letting LLMs do what they are good at — generating plausible solutions — and then employing program verification techniques to do what they are good at — verifying whether or not those solutions are correct.
6. Threats to Validity
Following Siegmund et al. (2015) and Feldt and Magazinius (2010) we identified a number of threats to the validity of this study. The closed nature of our experiments and the straightforward decision criteria (can Dafny compile a program? verify a program?) should ensure construct and internal validity: we are interested in the mechanistic question of whether a contemporary language model can generate code for a verifiable programming language, and we test that directly. Our subjective judgments were engaged primarily to classify mistakes and assess the fitness of the postconditions given the natural language prompts. Our qualitative evaluation used standard inspection methods to reach consensus, and all our data, prompts, coding, and intermediate analyses are available.
External validity considers how well these results would generalize to other language models or programming languages — or even to different versions of the tools we evaluated. In designing the study, we were aware that both the language model (GPT-4) and the programming language (Dafny) were evolving very rapidly: the version of GPT-4 we used had been available for only 14 days and Dafny 4.0 was released on 3rd March 2023; furthermore, these versions embodied significant improvements over previous versions of these tools. The only sure assertions we can make about the detailed results of this study are that they are unlikely to generalize: we expect the results will be different — most likely, better — if we repeated the study with later versions of the language model or the programming language. Unfortunately, there is no way to avoid these problems while studying contemporary software under continual, rapid, development: we consider the advantages of using the latest contemporary versions outweighs any increase in reproducibility we could get by e.g. using earlier stable versions that are several years old. We have, however, mitigated this threat as well as possible. We have identified the precise versions of the software and hardware that we used. The research artifacts of this study are publicly available at the companion repository.
7. Conclusion
We investigated the potential of large language models (GPT-4  and PaLM-2 ) to synthesize method specifications and bodies in the Dafny verification-aware programming language. We found that, given appropriate prompts, GPT-4  can generate Dafny methods, the pre- and postconditions needed to specify those methods, and the internal annotations required for Dafny to verify the methods against the specifications. Carefully engineered Chain of Thought (CoT) prompts with retrieval augmentation generated semantically similar few-shot examples explaining problem decomposition step by step yielded the best performance, generating the most post-conditions and minimizing unnecessary preconditions. Of the 178 problems in our test dataset, GPT-4 synthesized 103 methods with nontrivial specification that Dafny could verify: many of the remaining methods exhibited only minor errors. These findings underscore the importance of careful prompt design, and suggest that with further improvements in prompt design and more examples, GPT-4 can serve as the core of a “Programmer’s Verification Apprentice” for writing formally verified methods in Dafny. Although this work is exploratory, the results demonstrate great potential for more focused research on program synthesis with LLMs. We hypothesize that incorporating the tools and techniques of program verification may be an equally fruitful direction for research into generative models more generally, whenever their output must not only be plausible but also correct.
8. Data Availability
[…]